OpenStudy (anonymous):

Which shows the fractions in order from least to greatest? five-twelfths, two-ninths, five-sixths two-ninths, five-twelfths, five-sixths five-twelfths, two-ninths, five-sixths five-sixths, two-ninths, five-twelfths five-sixths, five-twelfths, two-ninths

OpenStudy (anonymous):

put them all over a common denominator, which would be a common multiple of 6, 9 and 12

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

We want to compare the fractions: \[\frac{ 5 }{ 12 } \frac{ 2 }{ 9 } \frac{ 5 }{ 6 } \] To start with, we need to find a common denominator. In this case we could use 3, adjusting each fraction to be over 3 by multiplying top and bottom by the same value. The new values are: \[\frac{ 5/4 }{ 3 } \frac{ 2/3 }{ 3 } \frac{ 5/2 }{ 3 } \] or \[\frac{ 1.25 }{ 3 } \frac{ .667 }{ 3 } \frac{ 2.5 }{ 3 } \] Now order from least to greatest: \[\frac{ .667 }{ 3 } \frac{ 1.25 }{ 3 } \frac{ 2.5 }{ 3 } \] Now you can just convert them back by looking above.

OpenStudy (anonymous):

There are a few ways to do it precisely, but with only three numbers separated by as much as these, you can estimate. Two ninths is less than a fourth. Five twelfths is about a half. Five sixths is almost one.

OpenStudy (anonymous):

\[\frac{ 5 }{12} , \frac{ 2 }{ 9 } , \frac{ 5 }{ 6 }\] common denominator is 36. So multiply them (by one) so that they are over 36. \[\frac{ 5 }{ 12 } \times \frac{ 3 }{ 3 } = \frac{ 15 }{ 36 }\] \[\frac{ 2 }{ 9 } \times \frac{ 4 }{ 4 } = \frac{ 8 }{ 36 }\] \[\frac{ 5 }{ 6 } \times \frac{ 6 }{ 6 } = \frac{ 30 }{ 36 }\] pretty much same as above.